# Towards Guaranteed Safe AI: A Framework for Ensuring Robust and Reliable AI Systems

### v2

@@ Abstract
Ensuring that AI systems reliably and robustly avoid harmful or dangerous behaviours is a crucial challenge, especially for AI systems with a high degree of autonomy and general intelligence, or systems used in safety-critical contexts. In this paper, we will introduce and define a family of approaches to AI safety, which we will refer to as guaranteed safe (GS) AI. The core feature of these approaches is that they aim to produce AI systems which are equipped with high-assurance quantitative safety guarantees. This is achieved by the interplay of three core components: a world model (which provides a mathematical description of how the AI system affects the outside world), a safety specification (which is a mathematical description of what effects are acceptable), and a verifier (which provides an auditable proof certificate that the AI satisfies the safety specification relative to the world model). We outline a number of approaches for creating each of these three core components, describe the main technical challenges, and suggest a number of potential solutions to them. We also argue for the necessity of this approach to AI safety, and for the inadequacy of the main alternative approaches.
@@ 1 Introduction
Critical infrastructure and safety-critical systems are required to comply with high safety standards. For example, aircrafts, nuclear power plants, and medical devices are subject to exceptionally rigorous safety certification. Moreover, it is plausible that there will soon be AI systems that are at least as safety-critical as these systems (e.g. as AI systems are increasingly deployed in safety-critical contexts, or given greater capabilities and autonomy), which means that they should be required to adhere to standards of safety that are at least as strict. Above some risk or capability thresholds, the burden of demonstrating such safety guarantees should be on the systems’ developers, and the provided evidence must be adequate to justify high confidence that the AI system is safe. We will argue that approaches based only on experimental tests are insufficient for producing such safety guarantees. We will also argue that GS AI presents research avenues that could plausibly produce such safety guarantees in a satisfactory and feasible manner, while remaining competitive with AI systems without such guarantees.
@@ 2.1 The AI Safety Problem
A number of prominent AI experts have argued that sufficiently advanced AI systems may threaten the survival of the human species, or lead to our permanent disempowerment, especially in the case of AI systems that are more intelligent than humans. Such concerns have been raised by Turing (1951) (who himself is refering to Butler, 1863), Russell (2019, 2024), Tegmark (2018), Bostrom (2014), Pearl (2019), Bengio (2023), Hinton (2023), Amodei (2023), and others. These experts have provided many different arguments in support of these concerns, which we will not reproduce here in full. However, here are a few very brief summaries of some of their central arguments:
@@ 3.1 Definition of GS AI
A high-assurance quantitative safety guarantee may take the form of a formal proof that the system always will adhere to some safety specification [^3] In this paper, we use the term “safety” not in the strict sense used in formal methods (Alpern & Schneider, 1987) but in the broader sense used in AI for the most important specifications of correct behaviour. (or distribution over safety specifications) for all inputs, relative to a model (or a distribution over models) of world dynamics. Alternatively, it may be a reliable and sound upper bound on the probability of violating a safety specification. However, especially when we cannot find a proof certificate, or cannot formally define the desirable or undesirable behaviour, it may also take the form of an estimate of an upper bound on the probability of harm, with the estimate asymptotically converging with computational resources in order to guarantee a constraint on desirable behaviour. This gives us the following definition (for a more formal definition, see Appendix A):
@@ 3.2 The World Model
The world model needs to answer queries about what would happen in the world as a result of a given output from the AI. It must also describe the state of the world at a level of granularity that is sufficient for expressing the safety specifications we are interested in. [^4] Note that a world model with a more abstract state-space may make it easier to express certain safety specifications, or make verification more tractable, but that this may come at the cost of making the predictions less accurate. Also note that the world model need not make predictions about arbitrary properties of the world, and even about properties on which they do make predictions, these predictions need not necessarily be precise. World models also serve to elucidate the AI system designers’ assumptions, and we must be mindful that those assumptions may hold only part of the time. Hence, the domain of applicability or epistemic uncertainty regarding different pieces of the world model must be represented and taken into account. For these reasons, the world model, or relevant aspects of it, should be auditable or monitorable at run time.
@@
  • Level 4: You use world models about real world phenomena that are formally verified as sound abstractions of fundamental physical laws.  [^5] By “sound abstraction” we mean the concept which has also been called “reverse simulation” or “oplax coalgebra morphism”. Formally, a sound abstraction of a dynamics $f:X\to\Delta X$ on a state space X is an abstracted state space A, an abstraction relation $a:X\to\mathcal{P}A$, and a dynamics $g:A\to\Delta A$ such that $f\fatsemi a\preceq a\fatsemi g$.
@@
Another approach is to learn a world model from data. However, this approach comes with a number of challenges. In particular, many machine learning methods are prone to being confidently incorrect in novel situations. On a theoretical level, this problem can be solved by Bayesian induction, which infers a distribution over world models (i.e. theories) from data (Kemp et al., 2010). Unfortunately, exact Bayesian induction is typically not computationally tractable (Hutter, 2003). However, recent advances in both Bayesian deep learning (Hollmann et al., 2023; Hu et al., 2023) and probabilistic programming suggest that approximate Bayesian induction can be made tractable (e.g.  Baydin et al., 2019a). Using deep learning to parameterize generative flow networks (Hu et al., 2023), one can train large neural networks to implicitly estimate and sample from the posterior distribution $P(T\mid D)$ over theories $T$ given data $D$, including cases where $T$ itself represents a structured (causal) model (Deleu et al., 2022; Ke et al., 2022; Deleu et al., 2023). Using probabilistic programming, one can leverage sequential Monte Carlo (SMC) methods to gradually infer a theory $T$ (represented as probabilistic program) as new data is observed (Saad et al., 2023), or by incrementally refining an initially coarse 3D world model (Gothoskar et al., 2023; Stuhlmüller et al., 2015). Both of these approaches can potentially be combined [^6] For example, the trained neural network can be used to guide SMC sampling (Zhao et al., 2024) and the resulting samples can in turn be used as training data for the neural network., and also benefit from more computation: by increasing the size or training time of the neural network used to approximate $P(T\mid D)$, convergence to the true posterior is expected. This means that we can continue training at run-time, or at least estimate the error made by the neural network through a sampling process. Similarly, by increasing the number of parallel samples, SMC methods asymptotically converge (Del Moral et al., 2006). We can further decrease the risks associated with insufficient training or sampling by encouraging the inferred theories to be human-inspectable (e.g. by conditioning on the constraint that theories can be converted to natural language and back with minimal error). If successful, this approach would produce a world model on Level 2 (or 4, if the model passes additional checks). See Bengio (2024) for a more complete discussion of this approach.
@@
A very ambitious approach would be to use world models that are formally verified as being sound abstractions of the basic laws of physics. Note that while physics is not a completed field, there is good reason to believe that our current best theories are completely accurate in certain domains: for example, quantum gravity effects appear irrelevant to contemporary AI systems (see e.g.  Carroll, 2021). We can therefore be confident that a system truly satisfies a specification if that specification is verified for that system relative to our best theories of physics, and that specification also includes the requirement that the system is not moved beyond the domains where our theories are known to be accurate. If successful, this would produce a world model on Level 4. One of the main challenges for this approach will be the immense computational complexity of producing and verifying such models. To reduce the computational complexity of this approach, approximate world models can be combined as compositional modules of one another, akin to how science approximates fluids as made of molecules made of atoms made of elementary particles, etc. For more details, see Tegmark & Omohundro (2023).
@@
Note that in many cases, it will likely be necessary for the world model to model the behaviour of humans. Moreover, humans are very complex, and it seems dubious to presuppose that it is possible to create a model of human behaviour that is both interpretable and highly accurate (especially noting that such a model itself would constitute an AGI system). There may thus in some cases be a fundamental trade-off between interpretability and predictive accuracy. However, note that interpretability can be maintained if human behaviour is modelled using nondeterminism. [^7] Here “nondeterminism” should be understood in the same sense as in “nondeterministic automata” and “nondeterministic Turing machines”, rather than as a synonym for “probabilistic”. In other words, the world model may specify a set or class of possible (potentially stochastic) behaviours for each human, and require the verifier to validate the safety specification for each behaviour profile in this set or class. The same applies to other highly complex systems. Also note that the levels in the classification above can be mixed within a single GS AI system. For example, a system can use stronger models of its engineering systems and weaker models of its interactions with the social domain (e.g., you may want to model your sensors with Level 5, but you will not be able to use that rigour for e.g. social phenomena).
@@ 3.3 The Safety Specification
Much of the AI safety literature formalises desirable system behavior in terms of reward functions (e.g.  Amodei et al., 2016). A safety specification is in general different from a reward function, though they include bounded reward functions as a special case. In particular, a safety specification may include properties defined by probabilistic temporal logics, causal counterfactual queries, or even hyperproperties, which reward functions cannot typically express (Seshia et al., 2018; Skalse & Abate, 2023a; Subramani et al., 2024). However, safety specifications cannot include unbounded evaluations of the world-state. [^9] The reason for this requirement is that it should be possible to completely satisfy the specification in order to relieve optimisation pressure. In order for it to be possible, it must at least be conceivable, i.e., the specification’s level of satisfaction needs to be bounded at least in theory. Safety specifications could be expressed in terms of PCTL (Hansson & Jonsson, 1994), or in terms of conjunction of linear inequalities of reachability probabilities, for example. Specification logics can also invoke neural components as predicates (Mao et al., 2019), allowing us to learn the semantics of concepts that are hard to manually specify.
The appropriate safety specification depends on the application domain and context (Khlaaf, 2023). Task-specific AI services often afford contextual operationalisations of safety, while more general-purpose AI assistants may require specifications that invoke broad definitions of harm (Beckers et al., 2022b; London & Heidari, 2023). For example, in the context of autonomous vehicles (or other embodied AI systems navigating in a human environment), the safety specification may contain rules about lawful and appropriate driving behaviour as well as specifications to avoid causing physical harm to others; for software system, the safety specification may seek to capture both its intended functionality and a lack of cyber vulnerabilities; or, in the context of automated medical research, the safety specification could bound the pathological potency of the identified molecules. Section 4 provides a number of concrete examples of GS AI, including their safety specifications.
@@
  • Level 1: The safety of the system is evaluated by a pool of human judges.
@@
Although specification can be challenging for highly competent and embodied AI systems, there are many low-hanging fruit tasks which do note require complex specification (or world modeling or verification) and which can provide great near-term value and incentive for further development. Examples include machines provably impossible to login to without correct credentials, DNA synthesizers that provably cannot synthesize certain pathogens, and AI hardware that is provably geofenced, time-limited (“mortal”) or equipped with a remote-operated throttle or kill-switch. Provably compliant sensors can be spefified to ensure “zeroization”, in which tampering with PCH is guaranteed to cause detection and erasure of private keys. For more details, see Tegmark & Omohundro (2023).
@@ 3.4 The Verifier
  • Level 10: Akin to level 9, but with the additional requirement that the proof is concise enough that humans can read, understand, and check it. 101010 For the formal proof case, level 9 suffices: Just as it is much harder to find a needle in a haystack than to verify that it is a needle, it is much harder to discover an algorithm with a compliance proof than to verify it — which can be done with just a few hundred lines of human-written code (Megill, 2023). This means that humans need not understand the system or proof, merely the the simple verifier code.
@@ 4 Examples of GS AI Solutions
Here we describe some concrete examples of practical problems and how they may be solved with a GS AI approach, which combine the different components described in the previous sections and fall along the different axes.
@@
Problem: Translate code in one programming language (e.g. C) to another (e.g. Rust) using AI systems (e.g. language models) with provable guarantees.
Specification: Functional correctness (the source program is equivalent to target program with respect to the relevant program state), and the generated target code does not have specified classes of security vulnerabilities.
@@ 5 Discussion
A third benefit of the GS research agenda is that it may produce AI safety solutions whose costs are amortised over time. Any potential safety measure (in AI or elsewhere) faces the issue that if said safety measure is costly to implement, then there is an incentive to disregard it. This reduces the value of approaches to AI safety that impose a high “safety tax”. For example, a comprehensive suite of rigorous empirical evaluations may be expensive and time-consuming to carry out, and this cost would (presumably) have to be paid again for each new AI system that is created. This would in turn create an incentive to save resources by cutting corners. By contrast, some approaches to GS AI may allow for most of these costs to be amortised, which may substantially reduce the incentive to disregard the corresponding safety measures. Once satisfactory safety specifications have been identified and scalable methods for formal verification have been developed, new AI systems could likely be verified against these safety predicates at a much lower marginal cost. Such solutions would also be more scalable in view of the impressive speed of AI progress that otherwise threatens to outpace progress in safety.
The range of applications for GS AI solutions, and with that also the scope of potential benefits, is large. A particularly significant type of application, which seems thus worth highlighting in the current context, is the securing of critical infrastructure such as water supply, power grids, nuclear plants, dams, information and communication technology, etc. Cybervulnerability of critical infrastructure may be one of the biggest threat vectors for catastrophic risks in the current world. For example, the US The Cyber Security & Infrastructure Security Agency (CISA) states that the “incapacitation or destruction [of critical infrastructure] would have a debilitating effect on security, national economic security, national public health or safety, or a combination thereof”. [^11] Retrieved from https://www.cisa.gov/topics/critical-infrastructure-security-and-resilience/critical-infrastructure-sectors on May 10, 2024. GS AI approaches have the potential to play a critical role in securing such critical infrastructure, thereby drastically bolstering civilisational resilience to both accidents and adversarial threats.
@@
Via the deployment infrastructure, we can modulate the time horizon at which the verified AI output is deployed. For example, at one level, the AI system may be deployed to operate autonomously with an infinite time horizon. To increase the reliability of safety guarantees, one may instead choose to deploy an AI system for a finite time horizon, such that, if at the end of that time horizon the AI system is not successfully verified for redeployment, operations shut down or transitions to a safe mode. The deployment systems can therefore also allow us to shorten the time horizon of outputs that need to be verified, thereby reducing the complexity of verification and increasing its tractability.

### v3

@@ Abstract
Ensuring that AI systems reliably and robustly avoid harmful or dangerous behaviours is a crucial challenge, especially for AI systems with a high degree of autonomy and general intelligence, or systems used in safety-critical contexts. In this position paper, we will introduce and define a family of approaches to AI safety, which we will refer to as guaranteed safe (GS) AI. The core feature of these approaches is that they aim to produce AI systems which are equipped with high-assurance quantitative safety guarantees. This is achieved by the interplay of three core components: a world model (which provides a mathematical description of how the AI system affects the outside world in a way that appropriately handles both Bayesian and Knighting uncertainty), a safety specification (which is a mathematical description of what effects are acceptable), and a verifier (which provides an auditable proof certificate that the AI satisfies the safety specification relative to the world model). We outline a number of approaches for creating each of these three core components, describe the main technical challenges, and suggest a number of potential solutions to them. We also argue for the necessity of this approach to AI safety, and for the inadequacy of the main alternative approaches.
@@ 1 Introduction
Critical infrastructure and safety-critical systems are required to comply with high safety standards. For example, aircraft, nuclear power plants, and medical devices are subject to exceptionally rigorous safety certification. Moreover, it is plausible that there will soon be AI systems that are at least as safety-critical as these systems (e.g. as AI systems are increasingly deployed in safety-critical contexts, or given greater capabilities and autonomy), which means that they should be required to adhere to standards of safety that are at least as strict. Above some risk or capability thresholds, the burden of demonstrating such safety guarantees should be on the systems’ developers, and the provided evidence must be adequate to justify high confidence that the AI system is safe. We will argue that approaches based only on experimental tests are insufficient for producing such safety guarantees. We will also argue that GS AI presents research avenues that could plausibly produce such safety guarantees in a satisfactory and feasible manner, while remaining competitive with AI systems without such guarantees.
@@ 2.1 The AI Safety Problem
A number of prominent AI experts have argued that sufficiently advanced AI systems may threaten the survival of the human species, or lead to our permanent disempowerment, especially in the case of AI systems that are more intelligent than humans. Such concerns have been raised by Butler (1863), Asimov (1942), Wiener (1950), Turing (1951), Good (1959), Minsky (1986), Moravec (1988), Vinge (1993), Joy (2000), Yudkowsky (2001), Bostrom (2002); as well as, more recently, Russell (2019, 2024), Tegmark (2018), Bostrom (2014), Pearl (2019), Bengio (2023), Hinton (2023), Amodei (2023), and others. These experts have provided many different arguments in support of these concerns, which we will not reproduce here in full. However, here are a few very brief summaries of some of their central arguments:
@@ 3.1 Definition of GS AI
A high-assurance quantitative safety guarantee may take the form of a formal proof that the system always will adhere to some safety specification [^3] In this paper, we use the term “safety” not in the strict sense used in formal methods (Alpern & Schneider, 1987) but in the broader sense used in AI for the most important specifications of correct behaviour. (or distribution over safety specifications) for all inputs, relative to a model (or a distribution over models) of world dynamics. Alternatively, it may be a reliable and sound upper bound on the probability of violating a safety specification. However, especially when we cannot find a proof certificate, or cannot formally define the desirable or undesirable behaviour, it may also take the form of an estimate of an upper bound on the probability of harm, with the estimate asymptotically converging with computational resources in order to guarantee a constraint on desirable behaviour. This gives us the following definition (for a more formal definition, see Appendix B):
@@ 3.2 The World Model
The world model needs to answer queries about what would happen in the world as a result of a given output from the AI. It must also describe the state of the world at a level of granularity that is sufficient for expressing the safety specifications we are interested in. [^4] Note that a world model with a more abstract state-space may make it easier to express certain safety specifications, or make verification more tractable, but that this may come at the cost of making the predictions less accurate. Also note that the world model need not make predictions about arbitrary properties of the world, and even about properties on which they do make predictions, these predictions need not necessarily be precise. A key challenge – which we will discuss more below – is finding a satisfactory solution to handling both Bayesian and Knightian uncertainty. World models also serve to elucidate the AI system designers’ assumptions, and we must be mindful that those assumptions may hold only part of the time. Hence, the domain of applicability and epistemic uncertainty regarding different pieces of the world model must be represented and taken into account. For these reasons, the world model, or relevant aspects of it, should be auditable or monitorable at run time.
@@
  • Level 4: You use world models about real world phenomena that are formally verified as sound abstractions of fundamental physical laws.  [^5] By “sound abstraction” we mean the concept which has also been called “reverse simulation” or “oplax coalgebra morphism”. Formally, a sound abstraction (of a dynamics $f:X\to\Delta X$ on a state space X) is an abstracted state space A, an abstraction relation $a:X\to\mathcal{P}A$, and a dynamics $g:A\to\Delta A$ such that $f\fatsemi a\preceq a\fatsemi g$.
@@
Another approach is to learn a world model from data. However, this approach comes with a number of challenges. In particular, many machine learning methods are prone to being confidently incorrect in novel situations. On a theoretical level, this problem can be solved by Bayesian induction, which infers a distribution over world models (i.e. theories) from data (Kemp et al., 2010). Unfortunately, exact Bayesian induction is typically not computationally tractable (Hutter, 2003). However, recent advances in both Bayesian deep learning (Hollmann et al., 2023; Hu et al., 2023) and probabilistic programming suggest that approximate Bayesian induction can be made tractable (e.g.  Baydin et al., 2019a). Using deep learning to parameterize generative flow networks (Hu et al., 2023), one can train large neural networks to implicitly estimate and sample from the posterior distribution $P(T\mid D)$ over theories $T$ given data $D$, including cases where $T$ itself represents a structured (causal) model (Deleu et al., 2022; Ke et al., 2022; Deleu et al., 2023). Using probabilistic programming, one can leverage sequential Monte Carlo (SMC) methods to gradually infer a theory $T$ (represented as probabilistic program) as new data is observed (Saad et al., 2023), or by incrementally refining an initially coarse 3D world model (Gothoskar et al., 2023; Stuhlmüller et al., 2015). Both of these approaches can potentially be combined [^6] For example, the trained neural network can be used to guide SMC sampling (Zhao et al., 2024) and the resulting samples can in turn be used as training data for the neural network., and also benefit from more computation: by increasing the size or training time of the neural network used to approximate $P(T\mid D)$, convergence to the true posterior is expected. This means that we can continue training at run-time, or at least estimate the error made by the neural network through a sampling process. Similarly, by increasing the number of parallel samples, sequential Monte Carlo (SMC) methods asymptotically converge (Del Moral et al., 2006). We can further decrease the risks associated with insufficient training or sampling by encouraging the inferred theories to be human-inspectable (e.g. by conditioning on the constraint that theories can be converted to natural language and back with minimal error). If successful, this approach would produce a world model on Level 2 (or 4, if the model passes additional checks). See Bengio (2024) for a more complete discussion of this approach.
@@
A very ambitious approach would be to use world models that are formally verified as being sound abstractions of the basic laws of physics. Note that while physics is not a completed field, there is good reason to believe that our current best theories are completely accurate in certain domains: for example, quantum gravity effects appear irrelevant to contemporary AI systems (see e.g.  Carroll, 2021). We can therefore be confident that a system truly satisfies a specification if that specification is verified for that system relative to our best theories of physics, and that specification also includes the requirement that the system is not moved beyond the domains where our theories are known to be accurate.  [^7] It is also worth acknowledging that even a perfect model of physical dynamics is insufficient for safety, since safety-critical queries (e.g. whether a given molecule is toxic to humans) presumably will depend on facts about the initial conditions (e.g. of human cells) that are not deducible from physics alone. This must be addressed by inference about initial conditions and boundary conditions from data and observations, tempered by appropriately conservative epistemic frameworks incorporating Bayesian and Knightian uncertainty. If successful, this would produce a world model on Level 4. One of the main challenges for this approach will be the immense computational complexity of producing and verifying such models. To reduce the computational complexity of this approach, approximate world models can be combined as compositional modules of one another, akin to how science approximates fluids as made of molecules made of atoms made of elementary particles, etc. For more details, see Tegmark & Omohundro (2023).
@@
Note that in many cases, it will likely be necessary for the world model to model the behaviour of humans. Moreover, humans are very complex, and it seems dubious to presuppose that it is possible to create a model of human behaviour that is both interpretable and highly accurate (especially noting that such a model itself would constitute an AGI system). There may thus in some cases be a fundamental trade-off between interpretability and predictive accuracy. However, note that interpretability can be maintained if human behaviour is modelled using nondeterminism. [^8] Here “nondeterminism” should be understood in the same sense as in “nondeterministic automata” and “nondeterministic Turing machines”, rather than as a synonym for “probabilistic”. In other words, the world model may specify a set or class of possible (potentially stochastic) behaviours for each human, and require the verifier to validate the safety specification for each behaviour profile in this set or class. Existing work on incorporating non-determinism into mathematical modelling comes form mathematics (Mio et al., 2021; Halpern & Leung, 2013; Kosoy, 2021; Liell-Cock & Staton, 2024), computer science (Brookes et al., 1984; Brookes & Roscoe, 1984; BenThacker et al., 2006; Zhang, 2022) as well as robotics LaValle & Hutchinson (1996). The same applies to other highly complex systems. Also note that the levels in the classification above can be mixed within a single GS AI system. For example, a system can use stronger models of its engineering systems and weaker models of its interactions with the social domain (e.g., you may want to model your sensors with Level 5, but you will not be able to use that rigour for e.g. social phenomena).
@@ 3.3 The Safety Specification
Much of the AI safety literature formalises desirable system behavior in terms of reward functions (e.g.  Amodei et al., 2016). A safety specification is in general different from a reward function, though they include bounded reward functions as a special case. In particular, a safety specification may include properties defined by probabilistic temporal logics, causal counterfactual queries, or even hyperproperties, which reward functions cannot typically express (Seshia et al., 2018; Skalse & Abate, 2023a; Subramani et al., 2024). However, safety specifications cannot include unbounded evaluations of the world-state. [^10] The reason for this requirement is that it should be possible to completely satisfy the specification in order to relieve optimisation pressure. In order for it to be possible, it must at least be conceivable, i.e., the specification’s level of satisfaction needs to be bounded at least in theory. Safety specifications could be expressed in terms of probabilistic computation tree logic (PCTL) (Hansson & Jonsson, 1994), or in terms of conjunction of linear inequalities of reachability probabilities, for example. Specification logics can also invoke neural components as predicates (Mao et al., 2019), allowing us to learn the semantics of concepts that are hard to manually specify.
The appropriate safety specification depends on the application domain and context (Khlaaf, 2023). Task-specific AI services often afford contextual operationalisations of safety, while more general-purpose AI assistants may require specifications that invoke broad definitions of harm (Beckers et al., 2022b; London & Heidari, 2023). [^11] See Appendix A for a short discussion of a potential definition of harm along the lines of Beckers et al. (2022a). For example, in the context of autonomous vehicles (or other embodied AI systems navigating in a human environment), the safety specification may contain rules about lawful and appropriate driving behaviour as well as specifications to avoid causing physical harm to others; for software systems, the safety specification may seek to capture both its intended functionality and a lack of cyber vulnerabilities; or, in the context of automated medical research, the safety specification could bound the pathological potency of the identified molecules. Section 4 provides a number of concrete examples of GS AI, including their safety specifications.
@@
  • Level 1: The safety of the system is evaluated by a pool of human judges based on their high-level intuitions and preferences.
@@
Although specification can be challenging for highly competent and embodied AI systems, there are many low-hanging fruit tasks which do not require complex specification (or world modeling or verification) and which can provide great near-term value and incentives for further development. Examples include machines provably impossible to login to without correct credentials, DNA synthesizers that provably cannot synthesize certain pathogens, and AI hardware that is provably geofenced, time-limited (“mortal”) or equipped with a remote-operated throttle or kill-switch. Provably compliant sensors can be spefified to ensure “zeroization”, in which tampering with PCH is guaranteed to cause detection and erasure of private keys. For more details, see Tegmark & Omohundro (2023).
@@ 3.4 The Verifier
  • Level 10: Akin to level 9, but with the additional requirement that the proof is concise enough that humans can read, understand, and check it. 121212 For the formal proof case, level 9 suffices: Just as it is much harder to find a needle in a haystack than to verify that it is a needle, it is much harder to discover an algorithm with a compliance proof than to verify it — which can be done with just a few hundred lines of human-written code (Megill, 2023). This means that humans need not understand the system or proof, merely the the simple verifier code.
@@ 4 Examples of GS AI Solutions
Here we describe some concrete examples of practical problems and how they may be solved with a GS AI approach, which combine the different components described in the previous sections and fall along the different axes. The examples range from current- or near-term tractable and economically viable applications, to ones that have not yet been realised due to their more ambitious scope.
@@
Problem: Translate code in one programming language (e.g. C) to another (e.g. Rust) using AI systems (e.g. language models) with provable guarantees. While some simple instances of this general problem have been solved (e.g., translating side effect-free functional programs from a source language to a domain-specific target language (Bhatia et al., 2024)), much remains to be done to solve the general code translation problem.
Specification: Functional correctness (the source program is equivalent to target program with respect to the relevant program state), and the generated target code does not have specified classes of security vulnerabilities. Defining functional correctness is easy in some cases (e.g., (Bhatia et al., 2024)) but more complicated in general since the source and target program structure can be very different and specifications can be partial.
@@ 5 Discussion
A third benefit of the GS research agenda is that it may produce AI safety solutions whose costs are amortised over time. Any potential safety measure (in AI or elsewhere) faces the issue that if said safety measure is costly to implement, then there is an incentive to disregard it. This reduces the value of approaches to AI safety that impose a high “safety tax”. For example, a comprehensive suite of rigorous empirical evaluations may be expensive and time-consuming to carry out, and this cost would (presumably) have to be paid again for each new AI system that is created. This would in turn create an incentive to save resources by cutting corners. By contrast, some approaches to GS AI may allow for most of these costs to be amortised, which may substantially reduce the incentive to disregard the corresponding safety measures. Once satisfactory safety specifications have been identified and scalable methods for formal verification have been developed, new AI systems could likely be verified against these specifications at a much lower marginal cost. Such solutions would also be more scalable in view of the impressive speed of AI progress that otherwise threatens to outpace progress in safety.
The range of applications for GS AI solutions, and with that also the scope of potential benefits, is large. A particularly significant type of application, which seems thus worth highlighting in the current context, is the securing of critical infrastructure such as water supply, power grids, nuclear plants, dams, information and communication technology, etc. Cybervulnerability of critical infrastructure may be one of the biggest threat vectors for catastrophic risks in the current world. For example, the US Cyber Security and Infrastructure Security Agency (CISA) states that the “incapacitation or destruction [of critical infrastructure] would have a debilitating effect on security, national economic security, national public health or safety, or a combination thereof”. [^13] Retrieved from https://www.cisa.gov/topics/critical-infrastructure-security-and-resilience/critical-infrastructure-sectors on May 10, 2024. GS AI approaches have the potential to play a critical role in securing such critical infrastructure, thereby drastically bolstering civilisational resilience to both accidents and adversarial threats.
@@
Via the deployment infrastructure, we can modulate the time horizon at which the verified AI output is deployed. For example, at one level, the AI system may be deployed to operate autonomously with an infinite time horizon. To increase the reliability of safety guarantees, one may instead choose to deploy an AI system for a finite time horizon, such that, if at the end of that time horizon the AI system is not successfully verified for redeployment, operations shut down or transition to a safe mode. The deployment systems can therefore also allow us to shorten the time horizon of outputs that need to be verified, thereby reducing the complexity of verification and increasing its tractability.